OpenStudy (anonymous):

***ANSWER FOR A MEDAL AND FAN!!!*** Write the equation in standard form. 2x^2+2y^2+12x-8y+24=0

OpenStudy (raden):

2x^2+2y^2+12x-8y+24=0 first, all divide by 2 x^2+y^2+6x-4y+12=0 then add by 1 to both sides, x^2+y^2+6x-4y+12+1=0+1 (x^2+6x+9)+(y^2-4y+4)=0+1 (x+3)^2 + (y-2)^2 = 1 it is an equation of circle with the centre (-3,2) and its radius r = 1